29. Solve. 6x + 4(3x – 2) = 6x + 2

Answer by adamchapman(301)

6x + 4(3x – 2) = 6x + 2
6x + 12x – 8 = 6x + 2
12x - 8 = 2
12x = 10
x = 1.2
I hope this helps,
Adam